OpenStudy (anonymous):

A child is given a push on a swing.Each time she swings back a little less, according to the equation d=2(0.85)^n , where d is the horizontal distance, in metre, from her position after n complete swings. After how many swings will the distance from rest be less than 15cm.

zepdrix (zepdrix):

Ooo this is annoying.. so they're measuring d in meters. And they want to know when d<15cm. I hate these unit changes >:c So they're really asking when d<.15meters, right? Did I do that conversion correctly?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

So we let our distance from the horizontal be .15 meters, \[\Large\rm .15=2(.85)^n\]And we need to solve for n, yes? Looks like we'll need to use logs.

zepdrix (zepdrix):

Divide each side by 2,\[\Large\rm .075=.85^n\]Then take the natural log of each side,\[\Large\rm \ln .075= \ln(.85^n)\] The reason we're using logs issssss, they help us get variables out of the exponent position.

OpenStudy (anonymous):

yaa you are right

zepdrix (zepdrix):

Rule of logs:\[\Large\rm \color{royalblue}{\log(a^b)=b \log (a)}\]Do you see how we can apply this rule to our problem here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Log (.85)^n = blog

OpenStudy (anonymous):

(.075)

zepdrix (zepdrix):

Ok great! The n comes out front as a multiplier. \[\Large\rm \ln .075= n ~\ln.85\] Understand how to solve for n from that point?

OpenStudy (anonymous):

yes i understand thanks a lot

zepdrix (zepdrix):

cool